# Learning to Synthesize Fault-Tolerant Quantum Circuits by Diagonalization

### v1

@@ VII Conclusion
We have demonstrated a novel approach to high precision unitary synthesis using discrete gate sets. By diagonalizing unitaries instead of directly inverting them, continuous degrees of freedom can be avoided then handled efficiently using mathematical decomposition methods. This enables us to handle more unitaries than other search-based synthesis methods which directly invert continuous rotations which is intractable for high levels of precision. We developed novel neural architectural components: global phase invariance and unitary perturbations, which enable deep learning entities to more effectively learn on unitary datasets. The effectiveness of our Reinforcement Learning based diagonalizing approach is demonstrated by synthesizing unitaries taken from a suite of partitioned quantum algorithms. Compared to the Quantum Shannon Decomposition, our approach reduces the number of non-trivial rotations (and therefore T-gates) in two-qubit unitaries by $82.8\%$ and $94.8\%$ in three-qubit unitaries. Because the diagonalizing approach is fast and able to find low gate count implementations of meaningful unitaries, it remains a promising technique for use in future fault-tolerant quantum compilers.
VIII Appendix
VIII-A Diagonal Distance
@@
Our goal is to determine when a state $s_{t}$ can be nearly inverted (satisfy the distance condition in Equation 6) by a completely diagonal unitary. Assume the state of the synthesis process is $s_{t}$, and it is “nearly diagonal.” This means that each row of $s_{t}$ is of the form
@@
VIII-B Global Phase Invariance

### v2

@@ 7. Conclusion
We also developed novel operations which can be applied to neural network architectures: global phase invariance and unitary perturbations. These modifications enable deep learning to more effectively learn on datasets of unitary matrices. The effectiveness of our Diagonalizing Value Network for Unitaries (DVNU) Reinforcement-Learning-based approach is demonstrated by synthesizing unitaries taken from a suite of partitioned quantum algorithms to very high precision. In this regime, only resource inefficient analytical methods are also able to find solutions. Compared to the Quantum Shannon Decomposition, our approach reduces the number of expensive $R_{Z}$ rotations (and therefore T-gates) by $83.5\%$ in two-qubit unitaries and $95.1\%$ in three-qubit unitaries. We demonstrate how our approach can be used to transpile future-term algorithms to fault-tolerant gate sets with very low approximation error. Using DVNU for transpilation results in up to a 16.8% reduction in non-Clifford gates for these algorithms. Our diagonalizing approach is fast and able to find low gate count implementations of meaningful unitaries, making it a promising technique for use in future fault-tolerant quantum compilers.
Acknowledgements.
The authors thank Sammuel Pfrommer for useful discussion about NeRF encodings. This work was supported by the DOE under contract DE-5AC02-05CH11231 through the Office of Advanced Scientific Computing Research (ASCR) Quantum Algorithms Team and Accelerated Research in Quantum Computing programs. This work was also supported by the NSF Challenge Institute for Quantum Computation (CIQC) program under award OMA-2016245 and through a CIQC seed funding award.
Appendix A Appendix
A.1. Diagonal Distance
@@
Our goal is to determine when a state $s_{t}$ can be nearly inverted (satisfy the distance condition in Equation 6) by a completely diagonal unitary. Say the state of the synthesis process is $s_{t}$, and it is “nearly diagonal.” This means that each row of $s_{t}$ is of the form
@@
A.2. Global Phase Invariance